Calculate the monthly finance charge for the following credit card transaction. Assume that it takes 10 days for a payment to be received and recorded and that the month is 30 days long. Round your answer to the nearesr cent. 3,000 balance, 21% rate, 150 payment, average daily method
Answer by Theo(13342)

365 days year is assumed.
the particular month in question is 30 days long.
your balance at the start of the month is 3,000.
the annual interest rate is 21%.
the payment is 150.00.
the interest rate you pay on the average daily balance depends on when in the month you send the payment in.
i'll make 2 assumptions.
first assumption is you send the payment in on the first of the month.
second assumption is you send thepayment in on the 20th day of the month.
here's how it works from what i could see from the following reference:
http://credit.about.com/od/creditcardbasics/qt/avgdailybalance.htm

YOU SEND THE PAYMENT IN ON THE FIRST DAY OF THE MONTH.
YOUR ACCOUNT IS CREDITED ON THE 11TH DAY OF THE MONTH.

average daily balance is 3000 from day 1 until day 10 (10 days)
average daily balance is 2850 from day 11 until day 30 (20 days)
average daily balance for the month is 10 * 3000 + 20 * 2850 = 87000 / 30 = 2900.
interest charge for the month is .21 * 2900 * 30 / 365 = 50.05 rounded to the nearest penny.

YOU SEND THE PAYMENT IN ON THE TWENTIETH DAY OF THE MONTH.
YOUR ACCOUNT IS CREDITED ON THE 30TH DAY OF THE MONTH.

average daily balance is 3000 from day 1 until day 29 (29 days)
average daily balance is 2850 from day 30 until day 30 (1 day)
average daily balance for the month is 3000 * 29 + 1 * 2850 = 89850 / 30 = 2995.
interest charge for the month is .21 * 2995 * 30 / 365 = 51.69 rounded to the nearest penny.

i could very easily be off by a day or 2 in my calculations depending on how the banks really do it, but, if the reference accurately describes it, then the procedure should be similar to what i just showed you.

here's another reference that says pretty much the same thing.
http://creditcardforum.com/blog/how-to-calculate-average-daily-balance-on-credit-card/

this method assumes interest is not compounded daily.
if interest is compounded daily, then you might wind up owing a little more because the interest earned each day is added to the principal so the principal for the next day contains a little more than it did the previous day.

bottom line is, if you're dealing with a particular bank, it's best to go to that bank and find out what their policy on charging interest is.

if it's just a how does that work type thing, these answers should be sufficient unless you were thinking of something different entirely.